OpenStudy (anonymous):

what is the slope intercept form for y+4=-7/5(x-5)

OpenStudy (anonymous):

So first you need to distribute the -7/5 on the right side. It will be: y+4=-7/5x+7. Then subtract 4 from both sides. y=-7/5x-3 That's the answer.

OpenStudy (texaschic101):

nicely done zeinas :)

OpenStudy (anonymous):

Thanks :) @texaschic101

OpenStudy (texaschic101):

just telling the truth

OpenStudy (shamil98):

@zeinas y+4=-7/5x+7. Then subtract 4 from both sides. y = -7/5x + 3 not -3